# A Survey on Statistical Theory of Deep Learning: Approximation, Training Dynamics, and Generative Models

### v1

@@ 2.1 Preliminary Problem Setting
Regression Task. Let $\mathcal{X}$ and $\mathcal{Y}\subset\mathbb{R}$ be the measurable feature space and output space. We denote $\rho$ as a joint probability measure on the product space $\mathcal{Z}:=\mathcal{X}\times\mathcal{Y}$, and let $\rho_{\mathcal{X}}$ be the marginal distribution of the feature space $\mathcal{X}$. We assume that the noisy data set $\mathcal{D}:=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{n}$ are generated from the non-parametric regression model
@@
Note the function estimator is taken from the feedforward neural network hypothesis space $\mathcal{F}(L,\mathbf{p},\mathcal{N})$ [^1] Henceforth, we will use a shorthand notation of $\mathcal{F}(L,\mathbf{p},\mathcal{N})$ as $\mathcal{F}$. Dependence on $(L,\mathbf{p},\mathcal{N})$ should be implicitly understood. defined in (2), and we denote the empirical minimizer of (4) as $\widehat{f}_{n}$. The goal is to estimate the regression function $f_{\rho}$ via feed-forward network $f\in\mathcal{F}$. Here, $f_{\rho}(\mathbf{x})$ is frequently assumed to be in certain function classes such as Hölder (i.e., $\mathcal{H}^{r}(\mathcal{X},K)$) or Sobolev (i.e., $W_{r}^{p}(\mathcal{X})$ for $1\leq p\leq\infty$) spaces. The notation $r(>0)$ represents the smoothness of functions, $K$ denotes the radius of Hölder ball, and $p$ denotes the $L_{p}$-norm used to measure magnitude of function derivatives. For more rigorous definitions of function classes, readers can refer Schmidt-Hieber (2020); Yarotsky (2017).
Decomposition of Excess risk. The excess risk is an important statistical object measuring the distance between the ground-truth function $f_{\rho}$ and function estimator via neural networks $\widehat{f}_{n}\in\mathcal{F}$. Mathematically, it can be shown that the excess risk is a difference between population risks of $f_{\rho}$ and $\widehat{f}_{n}$. We will be going to review the papers that utilize the following decomposition of the excess risk:

### v3

@@ 2.2 Statistical guarantees for regression tasks
Let $\mathcal{X}$ and $\mathcal{Y}\subset\mathbb{R}$ be the measurable feature space and output space. We denote $\rho$ as a joint probability measure on the product space $\mathcal{X}\times\mathcal{Y}$, and let $\rho_{\mathcal{X}}$ be the marginal distribution of the feature space $\mathcal{X}$. We assume that the noisy data set $\mathcal{D}:=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{n}$ are generated from the non-parametric regression model:
@@
Note that the papers introduced in this subsection always assume the empirical risk minimizer $\widehat{f}_{n}$ is obtainable, ignoring the optimization process. The function estimator $f$ is structurally regularized by $\mathcal{N}$ in $(L,\mathbf{p},\mathcal{N})$, which will be specified in sequel.
Under this setting, the excess risk is an important statistical object measuring the generalizability of the function estimator $\widehat{f}_{n}$ for unseen data in $\mathcal{X}$. Mathematically, it can be shown that it is a difference between population risks of $f_{\rho}$ and $\widehat{f}_{n}$ (See Chapter $13$ in Wainwright (2019)), which is $\mathbb{E}_{\mathbf{X}\sim\rho_{\mathcal{X}}}\big{[}(\widehat{f}_{n}(\mathbf{X})-f_{\rho}(\mathbf{X}))^{2}\big{]}$. The excess risk can be further decomposed as follows (Proposition $4.2$ in Suh et al. (2022)) :